[Ncep.nhc.nco_contacts] TSB Morning Rounds - Mon May 08, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Mon May 8 17:15:57 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig

----------

TAFB:
1. The Shiner N-AWIPS workstation was sluggish both on Saturday and Sunday,  
so TAFB forecasters rebooted it each time.

2. Problems occurred on Shiner when forecasters tried to overlay  
scatterometer data on satellite imagery via the "Data" item in the GUI  
menu. Forecasters are reminded that there is a known problem in N-AWIPS  
caused by the upgrade to the RHEL6 operating system: When calling up Data,  
the Accept and Close buttons do not appear when loading the Misc data type.  
Clicking on any other data type, such as "GRID" or "VGF", then back  
to "Misc", corrects the problem.

3. The clock on all N-AWIPS computers are slow by 1 minute. A NCO Help Desk  
ticket has been submitted requesting that all N-AWIPS workstations be  
connected to a NTP (Network Time Protocol) server for time synchronization.

4. Ongoing problems with the Cumulus N-AWIPS workstation: Cumulus is still  
underperforming. Enhanced logging was activated, which captured a 15-minute  
system freeze at 2111Z on Wednesday, May 3rd for analysis by NCO.
HSU:
1. ATCF issues: Hurricane Specialists received the red "guidance not ready"  
banner for EP90 at 00Z. Retrieved guidance and did not get SHIPS/LGEM the  
first time. Submitted guidance again at 0021Z and it was ready after 76  
seconds.

The issue with the slowness of the 00Z SHIPS/LGEM run of EP90 was because  
the grid decoding was not completed before the wallclock time limit  
expired. When the guidance was re-submitted, the decoding had completed.  
This sometimes happens when the guidance is submitted very early or the GFS  
model finishes running a bit late.

2. The 18Z HWRF for EP90 only returned a 0-hour tracker, so no HWFI was  
available for 00Z. Hurricane Specialists noticed that all of the SHIPS RII  
values at 18Z and 00Z for all thresholds were 0%, which seemed odd.  
The "%contribution" value for all but one of the predictors was 999.0,  
which might indicate something.

The problem with the missing rapid intensification index for the EP90  
invest was caused by the very low initial intensity (20 kt). There is a  
check for the input variables to see if they are within the bounds of the  
values in the developmental sample. If any predictor is outside of those  
bounds, %contribution is set to zero for that variable, and all others are  
set to an error flag, 999%. For the EP, there was never an RI case with an  
initial intensity of 20 kt, which triggered the error flag.
Other items:

NHC RFCs:
1. RFC 3206 - N-AWIPS v7.4.0 - Bugs fixes and table updates for N-AWIPS.
Includes:
- Updated NHC cone radii for 2017;
- NHC updates for 2017 (bug fixes, seasonal updates);
- Fixed missing 2nd tier marine zone; SFMAP performance improvement;
- Fixed PROB2CAT bugs that occurs while composing convective outlook and  
when lines are drawn outside the international boundaries;
- Bug fix in the ensemble diagnostic function (ENS_CVAL);
- and other changes.
Was to be implemented on April 27 at 1200Z; rescheduled for May 2.  
Postponed due to an issue that was found with the build; new implementation  
date is now Monday, May 8 (OPC - 1200Z; WPC, NCO, and NESDIS - 1400Z; CPC -  
2100Z).

2. RFC 3242 - Modify DBNet on WCOSS to begin sending some P-Surge files to  
IDP for NHC's new COOP machine. Was to be implemented on May 4 at 1630Z;  
postponed until Monday, May 8 at 1630Z due to a Critical Weather Day (CWD).

3. RFC 3255 - Turn on the ETSS v2.2 and PETSS v1 (Probabilistic  
Extratropical Storm Surge) parallel dataflows before the start of the  
30-day technical evaluation. Changes will take place on WCOSS and on NCODF  
in Boulder only. To be implemented on Tuesday, May 9 at 1430Z.

4. RFC 3189 - Make changes to IP tables on NHC systems on the 112 VLAN. A  
NHC user has requested SSH Access from the nhcgis.ncep.noaa.gov and  
nhcgisdev.ncep.noaa.gov systems to the 176 VLAN. Was to be implemented on  
April 18; postponed due to issues associated with RFC 3039 (router/switch  
upgrade). Implementation began on April 28 at 1600Z and should be completed  
by 1800Z on Tuesday, May 9.

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170508/472b69dd/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list